# The efficiency of black hole formation via collisions in stellar systems:

### v1

@@ 2 Methodology
We employ and test here the concept of the critical mass ($M_{\mathrm{crit}}$) introduced by Vergara et al. (2023), which is presented in subsection 2.1 accounting for the potential presence of an external potential, i.e. due to the presence of gas in the stellar system. For testing this concept, we gathered data from multiple sources of N-body simulations from which we determined both $M_{\mathrm{crit}}$ and the black hole formation efficiency ($\epsilon_{\mathrm{BH}}$), which is presented in subsection 2.2. These were computed using various initial conditions such as the initial stellar system mass $M_{\mathrm{i}}$, number of stars $N$, the half-mass radius $R_{\mathrm{h}}$, the stellar mass $m_{*}$, and radius $r_{*}$, besides the mass of the most massive object $M_{\mathrm{BH}}$ and age $\tau$ at the end of the simulations. A summary of the parameter space explored by different authors is provided in our Table 1 for simulations without external potential and in Table 2 for simulations with external potential. We further employ observations of NSCs, GCs, and UCDs (Table 3,  4, and  5, respectively). From these data, we use the current age $\tau$, mass $M_{\mathrm{i}}$, and effective radius $R_{\mathrm{eff}}$ of the stellar system, besides the mass of the black hole $M_{\mathrm{BH}}$.
@@ 2.1 Critical mass
The critical mass, introduced in Vergara et al. (2023) is
@@
In the presence of an external potential due to the presence of gas with a mass $M_{\mathrm{g}}$ within the stellar system, it is straightforward to generalize this expression under the assumption of virialization with the formalism employed by Reinoso et al. (2020), i.e.
@@ 2.2 Black hole formation efficiency
The other central quantity in this paper is the black hole formation efficiency ($\epsilon_{\mathrm{BH}}$), which is defined as the mass of the most massive object formed via collisions divided by the final stellar mass of the stellar system
@@
We note in particular here that the gas mass is not being considered in the definition of $\epsilon_{\mathrm{BH}}$, particularly as we are not considering the effects of accretion, and possible effects of the gas are only included through its effect on the steepening of the gravitational potential. A more detailed treatment of the gas would lead to accretion and would make the system more dissipative overall, thereby increasing the masses of the most massive objects to form, though such a treatment is beyond the scope of the present paper.
@@ 2.3 Numeric models
In this subsection, we compile data from different sources of N-body simulations. These simulations were performed using different codes, including bridge (Fujii et al., 2007), starlab (McMillan, 1996), nbody6 (Aarseth, 1999b, a), nbody6++gpu (Wang et al., 2015), and bifrost (Rantala et al., 2023), and have included different models for the star distribution, including profiles corresponding to Plummer (1911); King (1966); Miyamoto & Nagai (1975). They also employ different assumptions concerning the stellar mass function, including models with equal mass stars or employing an initial mass function (IMF) (Salpeter, 1955; Scalo, 1986; Kroupa, 2001). In the following, we will briefly describe the different simulations considered in this investigation.
@@
Fujii & Portegies Zwart (2013) explore the role of stellar collisions in virialized isolated cluster models, using a Salpeter (1955) IMF and the King (1966) density profile. While Sakurai et al. (2017) explore runaway collisions in first star clusters, using a Salpeter (1955) IMF. The distribution of stars was obtained from cosmological hydrodynamical simulations. Both studies use the bridge code (Fujii et al., 2007)
@@
Portegies Zwart et al. (1999) investigated runaway collisions in young compact star clusters, utilizing a Scalo (1986) IMF while Mapelli (2016) explored the impact of metallicity on runaway collisions, employing a Kroupa (2001) IMF. Both investigations employed a King (1966) profile and the starlab software (McMillan, 1996).
@@
Katz et al. (2015) investigated collisions in primordial star clusters at high redshift, using a Salpeter (1955) IMF. Reinoso et al. (2018) studied collisions in primordial star clusters, employing equal-mass stars. Reinoso et al. (2020) explored runaway collisions in dense star clusters, considering the effects of an external potential. Katz et al. (2015); Reinoso et al. (2018, 2020) use Plummer models (Plummer, 1911). Vergara et al. (2021) investigated collisions in flat and rotating clusters using a Miyamoto-Nagai distribution (Miyamoto & Nagai, 1975). All studies used the nbody6 code (Aarseth, 2000).
@@
Panamarev et al. (2019) simulated the Galactic center of the Milky Way using one million particles, despite the galactic center having approximately two orders of magnitude more stars, the computational limitations prevent the inclusion of a higher number of stars, the study employed a Kroupa (2001) IMF. Vergara et al. (2023) explored runaway collisions in NSCs, using models with equal-mass stars. Arca Sedda et al. (2023a, b, c), investigated star cluster evolution with up to one million stars, these are some of the most realistic simulations (dragon-II), including a Kroupa (2001) IMF, primordial binaries, stellar evolution, and dynamics of compact objects. All studies use a Plummer distribution (Plummer, 1911). The simulations employed the nbody6++gpu code (Wang et al., 2015).
Rizzuto et al. (2023) simulate dense clusters of low mass stars, using a Kroupa (2001) IMF and including an initial central black hole. This work was made with the code bifrost (Rantala et al., 2023).
@@
Key parameters for computing critical mass and black hole formation efficiency are summarized in Table 1. Besides in Table 2 we show simulations of Reinoso et al. (2020), where we include $q=M_{\mathrm{g}}/M_{\mathrm{i}}$. The complete Table 1 and Table 2 are available online (see Appendix A for more details).
@@
All the previously mentioned simulations employ N-body codes to study the stellar dynamics. While they share some similarities, they also have differences. The principal algorithm shared between codes is the Hermite integrator scheme, while bridge uses a $6^{\text{th}}$ order scheme. The other three codes use a $4^{\text{th}}$ order scheme. nbody6 and nbody6++gpu are direct N-body codes and have several algorithms in common, such as the Kustaanheimo-Stiefel regularization to solve close encounters (Stiefel & Kustaanheimo, 1965), the chain regularization (Mikkola & Aarseth, 1990, 1993), and the Ahmad-Cohen neighbor scheme which spatially splits the hierarchy of the stars to speed up computational calculations (Ahmad & Cohen, 1973). Indeed, nbody6++gpu is the improved version of nbody6, thus nbody6++gpu is optimized with GPU-accelerated supercomputing, which speeds up the calculations. This code also incorporates algorithms for solving stellar dynamics and recipes for stellar evolution proposed by Hurley et al. (2000, 2002), implemented by Kamlah et al. (2022). On the other hand, the most recent version of starlab is also a direct N-body code, adopted the stellar evolution recipe by Hurley et al. (2000), however in the past the code used recipes from Portegies Zwart & Verbunt (1996) for stellar evolution, incorporating Eggleton & Tout (1989) mass loss equations and a Schaerer (1999) model for massive stars losing their hydrogen envelope. bridge is a tree-direct hybrid N-body code. The internal region of the cluster is accurately solved through direct integration using the Hermite scheme (Nitadori & Makino, 2008), while the external motion is computed employing the tree algorithm (Barnes & Hut, 1986; Makino, 2004). The separation between the tree and direct calculations is achieved through Hamiltonian splitting (Kinoshita et al., 1990; Wisdom & Holman, 1991). Finally, bifrost is a direct summation N-body code accelerated by GPU with a hierarchical formulation of $4^{\text{th}}$ order forward symplectic integrator, originating in frost (Rantala et al., 2021) (the precursor of bifrost) where the integrator was initially derived and tested. We note that simulations with the code petar (Wang et al., 2020) are excluded from this analysis, as the code currently includes no treatment of collisions. More information about algorithms of stellar dynamics is available in the review of computational methods for collisional stellar systems made by Spurzem & Kamlah (2023).
Table 1: Summary of the initial conditions for the different sets of simulations. The first and second columns are $M_{\mathrm{i}}$ and $R_{\mathrm{h}}$ which are the initial mass and half mass radius of the stellar system, respectively. The third column is the number of stars. The fourth column $M_{\mathrm{BH}}$ is the black hole mass. The fifth is the time of the simulation. The sixth column shows the codes (B: bridge, SL: starlab, NB6: nbody6, NB6+: nbody6++gpu, and BF: bifrost) and the last column show the references.
@@
Table 2: Same columns as in Table 1, adding an extra column $q$.
@@ 2.4 Observations
In this subsection, we summarize the observational data on different types of stellar systems we employ in this analysis, considering NSCs, UCDs, and GCs. We summarize the critical parameters, including the age of the cluster, the effective radii, the masses of the stars, and the black hole within in Table 3 (NSCs), Table 4 (GCs), Table 5 (UCDs). Note that for all stellar systems, we assume that the initial mass is the sum of the current stellar mass and the mass of the black hole [^1] In principle additional mass loss could have occured, as we also find in our simulations. We checked that very similar results are obtained if we apply the same correction factor (accounting for mass loss) as we did in the case of simulations. However, in real physical systems, the situation may be more complex; particularly in the case of NSCs, they could also gain mass both due to in situ star formation and mergers with GCs. We therefore prefer to stick with a minimal set of assumptions within the analysis..
Table 3: Observational properties of NSCs. The columns are (1) stellar system ID, columns: (2, 3) BHs mass and reference, (4, 5) NSCs mass and reference, (6, 7) NSCs effective radius and reference, (8, 9) age and reference. Note for stellar systems where the age is not defined we assume a typical age, they are denoted by (*).
@@
Table 4: Observational properties of GCs. The columns are the same as in table 3.
@@
Table 5: Observational properties of UCDs. The columns are the same as in table 3.
@@
We assume that the initial mass of the galactic center was the combined mass of the black hole and NSC (Vergara et al., 2023). Additionally, we consider the initial effective radius to be one-tenth of its current size, as suggested by Banerjee & Kroupa (2017). We assume that the stars have solar properties (i.e. $1\leavevmode\nobreak\ \rm{\mathrm{M_{\odot}}}$ and $1\leavevmode\nobreak\ \rm{R_{\odot}}$). If the NSCs ages are not available we assume a typical age of 888\,8Gyr.
For the GCs we made the same assumption for the initial mass as for NSC (i.e. $M_{\rm i}=M_{\rm\textsc{GC}}+M_{\rm BH}$). These stellar configurations are one of the oldest objects in the Universe. The effective radius of isolated GCs remains constant through several relaxation timescales (Spitzer & Thuan, 1972), then we assume that the current size mirrors the initial size. GCs are old stellar systems. If the age is not available we assume a typical age of 121212\,12Gyr.
We analyze UCDs by taking into account assumptions from both GCs and NSCs regarding their effective radius, for the initial mass we consider $M_{\rm i}=M_{\rm\textsc{UCD}}+M_{\rm BH}$, given the uncertain nature of their formation. Some theories propose that UCDs could be remnants of tidal disruptions in dwarf galaxies interacting with larger ones, resulting in the stripping of outer layers and the formation of NSCs (Bassino et al., 1994; Bekki & Freeman, 2003; Pfeffer & Baumgardt, 2013). Alternatively, there is a suggestion that UCDs may originate from the merger of multiple GCs (Kroupa, 1998; Fellhauer & Kroupa, 2002; Fellhauer et al., 2002; Urrutia Zapata et al., 2019). In our investigation of UCDs, we consider two scenarios. UCD1 assumes a fixed radius during evolution, incorporating an external potential with $80\%$ dark matter due to observed high mass-to-light ratio values (Baumgardt & Mieske, 2008). UCD2, aligning more with Fellhauer & Kroupa (2006), posits an absence of dark matter and envisions an initial expansion like NSCs. Determining the age of UCDs is a challenging task (Francis et al., 2012). For those stellar systems that do not present an age measurement, we assume a typical age in between GCs and NSCs of 101010\,10Gyr.
@@ 3 Data analysis and comparison
In this section, we use the available data on simulations and observations to systematically test the concept of the critical mass. For this purpose, we first briefly recall the main differences between the respective datasets. In the case of simulations, we recall they represent idealized isolated systems. In most cases in the literature, their initial conditions, including the initial stellar mass $M_{\mathrm{i}}$ have been reported, as well as the main outcomes, such as the mass $M_{\mathrm{BH}}$ of the massive object that forms. In the case of observational data, on the other hand, we know about the current state of the system, i.e. the final stellar mass $M_{\mathrm{f}}$ and the mass of the most massive object $M_{\mathrm{BH}}$. We also note that real observational systems may be more complex than the simulation and may include additional physical processes. In the case of NSCs, gas physics, star formation and mergers with GCs could play some role in their evolution. For GCs and UCDs, on the other hand, it is more plausible that the stellar systems might be evolving for very long timescales essentially in isolation and without active star formation.
The computation of the critical mass as defined via Eq. 2 in principle requires knowledge of the initial conditions and is relatively trivial in the case of numerical simulations, while in the case of observations some assumptions or estimates may be necessary to determine the critical mass scale. On the other hand, the efficiency parameter $\epsilon_{\mathrm{BH}}$ defined in Eq. 3 requires knowledge of the final stellar mass and the mass of the most massive object. This appears straightforward from an observational perspective, while in simulations the final stellar mass has not always been documented and may thus require an estimate. In the following subsections, we will first analyse the results from simulations and observations separately and subsequently attempt a comparison between simulation and observation results.

### v2

@@ 2 Methodology
Here, we employed and tested the concept of the critical mass ($M_{\mathrm{crit}}$) introduced by Vergara et al. (2023), which is described in Section 2.1. We accounted for the potential presence of an external potential, namely, one that is due to the presence of gas in the stellar system. To test this concept, we gathered data from multiple sources of N-body simulations, from which we determined both $M_{\mathrm{crit}}$ and the black hole formation efficiency ($\epsilon_{\mathrm{BH}}$), which is presented in Section 2.2. These were computed using various initial conditions such as the initial stellar system mass, $M_{\mathrm{i}}$, number of stars, $N$, the half-mass radius, $R_{\mathrm{h}}$, the stellar mass, $m_{*}$, and radius, $r_{*}$, besides the mass of the most massive object, $M_{\mathrm{BH}}$, and age, $\tau$, at the end of the simulations. A summary of the parameter space explored by different authors is provided in our Table 1 for simulations without any external potential and in Table 2 for simulations with external potential. We further employed observations of NSCs, GCs, and UCDs (Table 3,  4, and  5, respectively). From these data, we use the current age, $\tau$; mass, $M_{\mathrm{i}}$; and effective radius, $R_{\mathrm{eff}}$, of the stellar system, besides the mass of the black hole, $M_{\mathrm{BH}}$.
@@ 2.1 Critical mass
The critical mass, introduced in Vergara et al. (2023) is:
@@
In the presence of an external potential due to the presence of gas with a mass, $M_{\mathrm{g}}$, within the stellar system, it is straightforward to generalize this expression under the assumption of virialization with the formalism employed by Reinoso et al. (2020), namely,
@@ 2.2 Black hole formation efficiency
The other central quantity in this paper is the black hole formation efficiency ($\epsilon_{\mathrm{BH}}$), which is defined as the mass of the most massive object formed via collisions divided by the final stellar mass of the stellar system:
@@
We note, in particular, that the gas mass is not being considered in the definition of $\epsilon_{\mathrm{BH}}$, especially as we are not considering the effects of accretion, whereas the possible effects of the gas are only included through its effect on the steepening of the gravitational potential. A more detailed treatment of the gas would lead to accretion and would make the system more dissipative overall, thereby increasing the masses of the most massive objects to form. However, such a treatment is beyond the scope of the present paper.
@@ 2.3 Numeric models
In this subsection, we compile the data from different sources of N-body simulations. These simulations were performed using different codes, including bridge (Fujii et al., 2007), starlab (McMillan, 1996), nbody6 (Aarseth, 1999b, a), nbody6++gpu (Wang et al., 2015), and bifrost (Rantala et al., 2023). These codes have included different models for the star distribution, including profiles corresponding to Plummer (1911); King (1966); Miyamoto & Nagai (1975). They also employ different assumptions concerning the stellar mass function, including models with equal mass stars or employing an initial mass function (IMF) (Salpeter, 1955; Scalo, 1986; Kroupa, 2001). In the following, we briefly describe the different simulations considered in this investigation.
@@
Fujii & Portegies Zwart (2013) explored the role of stellar collisions in virialized isolated cluster models, using a Salpeter (1955) IMF and the King (1966) density profile. While Sakurai et al. (2017) explored runaway collisions in first star clusters, using the IMF from Salpeter (1955). The distribution of stars was obtained from cosmological hydrodynamical simulations. Both studies use the bridge code (Fujii et al., 2007)
@@
Portegies Zwart et al. (1999) investigated runaway collisions in young compact star clusters, utilizing a Scalo (1986) IMF, while Mapelli (2016) explored the impact of metallicity on runaway collisions, employing a Kroupa (2001) IMF. Both investigations employed a King (1966) profile and the starlab software (McMillan, 1996).
@@
Katz et al. (2015) investigated collisions in primordial star clusters at high redshift, using a Salpeter (1955) IMF. Reinoso et al. (2018) studied collisions in primordial star clusters, employing equal-mass stars. Reinoso et al. (2020) explored runaway collisions in dense star clusters, considering the effects of an external potential. Katz et al. (2015); Reinoso et al. (2018, 2020) used Plummer models (Plummer, 1911). Vergara et al. (2021) investigated collisions in flat and rotating clusters using a Miyamoto-Nagai distribution (Miyamoto & Nagai, 1975). All studies used the nbody6 code (Aarseth, 2000).
@@
Panamarev et al. (2019) simulated the Galactic center of the Milky Way using one million particles, despite the galactic center having approximately more stars by two orders of magnitude, the computational limitations prevent the inclusion of a higher number of stars, the study employed the IMF from Kroupa (2001). Vergara et al. (2023) explored runaway collisions in NSCs, using models with equal-mass stars. Arca Sedda et al. (2023a, b, c), investigated star cluster evolution with up to one million stars, these are some of the most realistic simulations (dragon-II), including a Kroupa (2001) IMF, primordial binaries, stellar evolution, and dynamics of compact objects. All studies use a Plummer distribution (Plummer, 1911). The simulations employed the nbody6++gpu code (Wang et al., 2015). Rizzuto et al. (2023) simulated dense clusters of low mass stars, using a Kroupa (2001) IMF and including an initial central black hole. This work was made with the bifrost code (Rantala et al., 2023).
@@
The key parameters for computing the critical mass and black hole formation efficiency are summarized in Table 1. In addition, in Table 2, we show the simulations of Reinoso et al. (2020), where we include $q=M_{\mathrm{g}}/M_{\mathrm{i}}$. The complete Tables 1 and 2 are available online (see Appendix A for more details).
@@
All the previously mentioned simulations employ N-body codes to study the stellar dynamics. While they share some similarities, they also have differences. The principal algorithm shared between codes is the Hermite integrator scheme, while bridge uses a sixth-order scheme. The other three codes use a fourth-order scheme. Moreover, nbody6 and nbody6++gpu are direct N-body codes and have several algorithms in common, such as the Kustaanheimo-Stiefel regularization to solve close encounters (Stiefel & Kustaanheimo, 1965), chain regularization (Mikkola & Aarseth, 1990, 1993), and the Ahmad-Cohen neighbor scheme which spatially splits the hierarchy of the stars to speed up computational calculations (Ahmad & Cohen, 1973). Indeed, nbody6++gpu is the improved version of nbody6, thus nbody6++gpu is optimized with GPU-accelerated supercomputing, which speeds up the calculations. This code also incorporates algorithms for solving stellar dynamics and recipes for stellar evolution proposed by Hurley et al. (2000, 2002), implemented by Kamlah et al. (2022). On the other hand, the most recent version of starlab is also a direct N-body code, adopted the stellar evolution recipe by Hurley et al. (2000); however, in the past, the code used recipes from Portegies Zwart & Verbunt (1996) for stellar evolution, incorporating the Eggleton & Tout (1989) mass-loss equations and the Schaerer (1999) model for massive stars losing their hydrogen envelope. bridge is a tree-direct hybrid N-body code. The internal region of the cluster is accurately solved through direct integration using the Hermite scheme (Nitadori & Makino, 2008), while the external motion is computed employing the tree algorithm (Barnes & Hut, 1986; Makino, 2004). The separation between the tree and direct calculations is achieved through Hamiltonian splitting (Kinoshita et al., 1990; Wisdom & Holman, 1991). Finally, bifrost is a direct summation N-body code accelerated by GPU with a hierarchical formulation of fourth-order forward symplectic integrator, originating in frost (Rantala et al., 2021), the precursor of bifrost, where the integrator was initially derived and tested. We note that the code petar (Wang et al., 2020) has been used to study mergers in star clusters (Wang et al., 2024; Barber et al., 2024) and population III star clusters (Wang et al., 2022; Liu et al., 2023). Although we did not include this data, Wang et al. (2022) simulations fall within $M_{i}/M_{crit}<0.001$, keeping our conclusions unchanged. More information about algorithms of stellar dynamics is available in the review of computational methods for collisional stellar systems made by Spurzem & Kamlah (2023).
Table 1: Initial conditions for simulation sets
@@
Table 2: Initial conditions for simulation sets with external potential
@@ 2.4 Observations
In this subsection, we summarize the observational data on different types of stellar systems we employ in this analysis, considering NSCs, UCDs, and GCs. We summarize the critical parameters, including the age of the cluster, the effective radii, the masses of the stars, and the black hole within in Table 3 (NSCs), Table 4 (GCs), and Table 5 (UCDs). We note that for all stellar systems, we assumed that the initial mass is the sum of the current stellar mass and the mass of the black hole [^3] In principle, additional mass loss could have occured, as we also find in our simulations. We checked that very similar results are obtained if we apply the same correction factor (accounting for mass loss) as we did in the case of simulations. However, in real physical systems, the situation may be more complex; particularly in the case of NSCs, they could also gain mass both due to in situ star formation and mergers with GCs. We therefore prefer to stick with a minimal set of assumptions within the analysis..
Table 3: Observational properties of NSCs.
@@
Table 4: Observational properties of GCs.
@@
Table 5: Observational properties of UCDs.
@@
We assumed that the initial mass of the galactic center was the combined mass of the black hole and NSC (Vergara et al., 2023). Additionally, we considered the initial effective radius to be one-tenth of its current size, as suggested by Banerjee & Kroupa (2017). We assumed that the stars have solar properties (i.e., $1\leavevmode\nobreak\ \rm{\mathrm{M_{\odot}}}$ and $1\leavevmode\nobreak\ \rm{R_{\odot}}$). If the NSCs ages were not available we assume a typical age of 888\,8Gyr.
For the GCs, we made the same assumption for the initial mass as for NSC (i.e., $M_{\rm i}=M_{\rm\textsc{GC}}+M_{\rm BH}$). These stellar configurations are one of the oldest objects in the Universe. The effective radius of isolated GCs remains constant through several relaxation timescales (Spitzer & Thuan, 1972), then we assume that the current size mirrors the initial size. GCs are old stellar systems. If the age was not available, we assumed a typical age of 121212\,12Gyr.
We analyzed the UCDs by taking into account assumptions from both GCs and NSCs regarding their effective radius; for the initial mass, we consider $M_{\rm i}=M_{\rm\textsc{UCD}}+M_{\rm BH}$, given the uncertain nature of their formation. Some theories propose that UCDs could be remnants of tidal disruptions in dwarf galaxies interacting with larger ones, resulting in the stripping of outer layers and the formation of NSCs (Bassino et al., 1994; Bekki & Freeman, 2003; Pfeffer & Baumgardt, 2013). Alternatively, there is a suggestion that UCDs may originate from the merger of multiple GCs (Kroupa, 1998; Fellhauer & Kroupa, 2002; Fellhauer et al., 2002; Urrutia Zapata et al., 2019). In our investigation of UCDs, we considered two scenarios. In particular, UCD1 assumes a fixed radius during evolution, incorporating an external potential with $80\%$ dark matter due to observed high mass-to-light ratio values (Baumgardt & Mieske, 2008); UCD2, aligning more with Fellhauer & Kroupa (2006), posits an absence of dark matter and envisions an initial expansion like NSCs. Determining the ages of UCDs is a challenging task (Francis et al., 2012). For those stellar systems that do not present an age measurement, we assumed a typical age in between GCs and NSCs of 101010\,10Gyr.
@@ 3 Data analysis and comparison
In this section, we use the available data on simulations and observations to systematically test the concept of the critical mass. For this purpose, we first briefly recall the main differences between the respective datasets. In the case of simulations, we recall they represent idealized isolated systems. In most cases in the literature, their initial conditions, including the initial stellar mass, $M_{\mathrm{i}}$, have been reported, as well as the main outcomes, such as the mass $M_{\mathrm{BH}}$ of the massive object that forms. In the case of observational data, on the other hand, we know about the current state of the system, namely, the final stellar mass, $M_{\mathrm{f}}$, and the mass of the most massive object, $M_{\mathrm{BH}}$. We also note that real observational systems may be more complex than the simulation and may include additional physical processes. In the case of NSCs, gas physics, star formation, and mergers with GCs could play some role in their evolution. For GCs and UCDs, on the other hand, it is more plausible that the stellar systems might be evolving for very long timescales essentially in isolation and without active star formation.
The computation of the critical mass (as defined via Eq. 2) does, in principle, require knowledge of the initial conditions and is relatively trivial in the case of numerical simulations, while in the case of observations some assumptions or estimates may be necessary to determine the critical mass scale. On the other hand, the efficiency parameter $\epsilon_{\mathrm{BH}}$ defined in Eq. 3 requires knowledge of the final stellar mass and the mass of the most massive object. This appears straightforward from an observational perspective, whereas in simulations, the final stellar mass has not always been documented and may thus require an estimate. In the following subsections, we first analyse the results from simulations and observations separately and subsequently attempt to make a comparison between simulation and observation results.